# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v3

@@ Appendix C Comparison with Lee & Jang (2023)
$\begin{split}2/\eta\ \geq\ \mathrm{IAS}\ &=\ \frac{1}{\|\nabla f\|^{2}}\left(\nabla f^{\top}\cdot\nabla^{2}f\cdot\nabla f\mathbb{E}\left[\zeta_{B}^{\top}\cdot\nabla^{2}f\cdot\zeta_{B}\right]\right)\\ &=\ \frac{1}{\|\nabla f\|^{2}}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f,v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B},v_{i}\rangle^{2}]\,\big{)}.\end{split}$ (13)
However, after $T$ steps we can lower bound the loss with an increasing sequence which eventually diverges. Moreover, the loss asymptotically will approach the sequence with a linear convergence rate. Precisely, define the noise $\zeta_{B}$ such that $\nabla f_{B}=\nabla f+\zeta_{B}$. Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of Lee & Jang (2023) we may have decrease in loss in one step, indeed
$\begin{split}f(x_{T})-f(x_{0})\ &=\ \sum_{t=0}^{T-1}\left(\frac{\eta^{2}}{2}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f(x_{t}),v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{i}\rangle^{2}]\,\big{)}-\|\nabla f(x_{t})\|^{2}\right)\\ &\geq\ \sum_{t=0}^{T-1}\underbrace{\left(\frac{\eta^{2}}{2}\lambda_{1}-\eta\right)\,\langle\nabla f(x_{t}),v_{1}\rangle^{2}}_{positive}\,+\,\underbrace{\frac{\eta^{2}}{2}\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{1}\rangle^{2}]}_{positive}.\end{split}$ (14)
Both the quantities under-braced above–the parallel component to $v_{1}$–explode with the time $t$. This implies, in particular, that imposing $IAS\leq 2/\eta$ may implies one-step stability, but not trajectory stability.
Our quantity, in turn, factors in the variability that the eigenspaces have on different batches. Our quantity is also strictly bigger, thus imposing it to be equal to $2/\eta$ induces a lower sharpness, see, e.g, Appendices G and I.
Appendix D Stability for Noise-Injected Gradients Descents
We will see here that what we showed above is the case of mini-batch algorithms but not of many other versions of stochastic algorithms considered in theoretical work or practice.
We deal here with the case of noisy gradients, which is the case usually analyzed in the practice of theory works, see Section 6. While in minibatch stochastic gradient descent (SGD), the step function is
$x_{t+1}=x_{t}-\eta\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x_{t}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
and the noise comes from mini batching as shown above, in noisy gradient descents we do not care of a possible way to rewrite $A,b,c$ as sum of other terms. Just, there exists a random variable $\zeta_{t}$, independent for different times $t_{1}\neq t_{2}$, such that the $t-th$ step is
$x_{t+1}=x_{t}-\eta\left(\mathbf{A}\mathbf{(x_{t}+\zeta_{t})}+b\right).$
This implies that
$x_{t+1}-x^{*}=\left(I+\eta\mathbf{A}\right)\big{(}x_{t}-x^{*}\big{)}\ +\ A\zeta_{t}.$
This imply that the Hessian does not change from step to step. The expected stability formulation is actually the one of Lee & Jang (2023) in the sense of Cohen et al. (2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
This, in particular, establishes mathematically and agrees with what we see empirically in Section 6.
Appendix E Sampling the Mini Batches
For simplicity let us work with the one dimensional case. The point in this section is that In the one dimensional case where $\mathbf{A}_{i}=\mathbf{a}_{i},\mathbf{b}_{i},\mathbf{c}_{i}\in\mathbb{R}$ for all $i\in\{1,2,\ldots,n\}$ we have that
$\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t+1}-x^{*}]=\underbrace{\mathbb{E}\left[\left(1-\eta\frac{1}{b}\sum_{i\in B}\mathbf{a}_{i}\right)\big{(}x_{t}-x^{*}\big{)}\right]}_{(1-\eta\mathbf{a})\cdot\mathbb{E}[x_{t+1}-x^{*}]\text{ when independent steps}}\ -\ \underbrace{\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}\left[\frac{1}{b}\sum_{i\in B}\mathbf{a}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right]}_{0\text{ when uniform sampling}}.$ (15)
When $a_{i},b_{i}$ are sampled uniformly within the dataset and independently of the step $t$, then this becomes
$\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t+1}-x^{*}]=(1-\eta\mathbf{a})\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t}-x^{*}]$ (16)
and the stability threshold is the same as for full-batch GD or any other momentum version, for instance.
In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of Beneventano (2023) as the tools previously developed by Mishchenko et al. (2020); Smith et al. (2021) only work for infinitesimal step size and the tools in Appendix E of Beneventano (2023) are the only ones which allow us to consider more products and big learning rate. Let $k=n/b$ be the number of batches in an epoch and call $B_{j}$ the batch of $j-th$ step and $\mathbf{a}_{B_{j}},\mathbf{b}_{B_{j}}$ the averages of $\mathbf{a}$ and $\mathbf{b}$ over the batch $B_{j}$.
$x_{k}-x^{*}=\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\big{(}x_{t}-x^{*}\big{)}\ -\ \sum_{j=1}^{k}\left[\prod_{h=j+1}^{k}\left(1-\eta\mathbf{a}_{B_{h}}\right)\right]\big{(}\mathbf{a}_{B_{j}}x^{*}+\mathbf{b}_{B_{j}}\big{)}.$ (17)
Next note that by taking the expectation of the term $\prod_{j=1}^{k}\left(1+\eta\mathbf{a}_{B_{j}}\right)$ we obtain
[…]

### v4

@@ Appendix B Quadratic Setting: Batch Sharpness and GNI
Note that the fact that $\mu^{\top}(H_{i}-H)\mu$ is centered in zero, implies that
[…]
This establish the first thesis of the lemma and implies that
$|Y^{\!\top}(H_{i}-H)Y|\;\leq\;\|H_{i}-H\|_{2}\,\left(\|\widetilde{Y}\|^{2}+2|\mu^{\top}\widetilde{Y}|+\lVert\mu\rVert^{2}\right).$
Taking the expectation we have
$|\Delta|\leq\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2\lVert\mu\rVert_{2}^{2}\cdot\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{2}]+\lVert\mu\rVert_{2}^{4}}.$
This concludes the proof of the lemma. ∎
$\Delta$ measures how curvature noise $(H_{i}-H)$ correlates with noise $Y$. Lemma 5 shows this interaction cannot exceed the root product of the moments of two natural quantities.
B.6 Main trace inequality
Proposition 3.
We have that Batch Sharpness is smaller than GNI if and only if
$\mathcal{B}C_{0}\;\geq\;\lVert\mu\rVert_{2}^{2}\cdot(\tilde{\Delta}+2\mathcal{D}).$ (21)
In particular,
$\mathcal{B}\,\mathcal{C}-\mathcal{A}\,\|\mu\|^{2}\;\geq\;\mathcal{B}C_{0}-\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}\lVert\mu\rVert^{2}.$ (22)
This implies that Batch Sharpness is smaller than GNI up to $O(\eta)$ at the stationary distribution when
$C_{0}\geq\frac{\eta}{2}\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}.$ (23)
Proof.
Note that $\mathcal{A}\lVert\mu\rVert^{2}\leq\mathcal{B}\lVert\mu\rVert^{2}+|\Delta|\lVert\mu\rVert^{2}$. Applying Lemma 5 and noticing that $\mathcal{C}=C_{0}+\lVert\mu\rVert^{2}$ concludes the proof of (22). Next note that up to $\mathcal{O}(\eta)$ we have that GNI $=\mathcal{B}/\lVert\mu\rVert^{2}=2/\eta$ at the stationary distribution. This proves the second part (23) and establishes Proposition 3. ∎
B.7 The moments of the oscillations
1. Centering first simplifies the ratio
Because $C_{0}$ is defined with the centred residual $\widetilde{Y}$, we compare
$\frac{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}{C_{0}^{2}}\;=\;\frac{\mathbb{E}[\|\widetilde{Y}\|^{4}]}{C_{0}^{2}}+2\,\frac{\|\mu\|^{2}}{C_{0}}+\frac{\|\mu\|^{4}}{C_{0}^{2}}.$
The first term is the kurtosis-type ratio $\kappa_{Y}\;:=\;\mathbb{E}[\|\widetilde{Y}\|^{4}]/C_{0}^{2}.$
2. Exact formula in the Gaussian quadratic setting
Assume:
  • $x_{i}\!\mid\!H_{i}\sim\mathcal{N}(\theta^{\star},\Sigma_{x})$ – an isotropic sampling model for simplicity;
  • $H_{i}$ is deterministic, i.e. $H_{i}\equiv H\succeq 0$.
Then $\xi:=Y=H(\theta^{\star}-x_{i})\sim\mathcal{N}\!\bigl{(}0,\;S\bigr{)}$ with $S:=H\Sigma_{x}H$ and $\widetilde{Y}\equiv Y$ because $\mu=0$. By Isserlis’ theorem
$\mathbb{E}[\|Y\|^{4}]\;=\;(\operatorname{tr}S)^{2}+2\operatorname{tr}(S^{2}),\qquad C_{0}\;=\;\operatorname{tr}S.$
Therefore
[…]
  • Isotropic case $S=\sigma^{2}I_{d}$Longrightarrow $\kappa=1+\tfrac{2}{d}$, e.g. $=3$ in $d=1$, $=1.4$ in $d=5$, $\to 1$ for large $d$.
  • Rank-one noise $S=\sigma^{2}uu^{\!\top}$ ($\|u\|=1$$\Longrightarrow$ $\kappa=3$ – the maximal value compatible with Gaussianity.
  • The interval bounds follow from $\operatorname{tr}(S^{2})\in[\tfrac{1}{d}(\operatorname{tr}S)^{2},\,(\operatorname{tr}S)^{2}]$.
3. Effect of random curvature $H_{i}$
If $H_{i}$ varies but is independent of $x_{i}$ and of $\theta$ a priori, the second moment becomes
$C_{0}\;=\;\mathbb{E}_{H}\!\bigl{[}\operatorname{tr}(H\Sigma_{x}H)\bigr{]},$
while
$\mathbb{E}[\|Y\|^{4}]\;=\;\mathbb{E}_{H}\Bigl{[}(\operatorname{tr}(H\Sigma_{x}H))^{2}+2\,\operatorname{tr}\!\bigl{(}(H\Sigma_{x}H)^{2}\bigr{)}\Bigr{]}.$
Define $a:=\mathbb{E}_{H}[\operatorname{tr}(H\Sigma_{x}H)],\;b:=\mathbb{E}_{H}[\operatorname{tr}(H\Sigma_{x}H)^{2}],\;c:=\mathbb{E}_{H}[\operatorname{tr}((H\Sigma_{x}H)^{2})].$ Then
[…]
B.8 Closing up
In the setting above, thus, we can transform (23) in the following inequality
$1\geq\frac{\eta}{2}\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\frac{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}{C_{0}^{2}}}$ (26)
and the previous subsection tells us that there exists a universal constant $\kappa_{Y}$ which is a form of kurtosis of our steps (which depends on our distribution) such that Batch Sharpness is smaller than GNI whenever
$\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\leq\;\frac{2}{\eta(\kappa_{Y}+\mathcal{O}(\eta))}.$ (27)
Proposition 4 (Stationary trace inequality).
For every $\eta$ such that $0<\eta<2/\lambda_{\max}$ we have
$\mathcal{A}\,\|\mu\|^{2}\;\leq\;\mathcal{B}\,\mathcal{C}$ (28)
up to $\mathcal{O}(\eta)$ when
$\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\leq\;\frac{2}{\eta\kappa_{Y}}+\mathcal{O}(\eta).$ (29)
In particular, at the beginning of the training for randomly initialized weights this is the case as the expectation of the LHS is $\mathcal{O}(\log(d))$ where $d$ is the number of parameters.
Proof.
The theorem follows from (27) and (25). ∎
Interpretation.
Inequality (28) says: “the loss under random curvature cannot align too strongly with the squared prediction bias.” When either (i) the stationary iterate is unbiased ($\mu=0$) or (ii) the curvature is deterministic, the bound is tight.
Appendix C Gradients explode above the EoSS: Proof of Theorem 1
We now compute the update of the norm of the gradients $\mathbb{E}_{i}[\lVert Y_{i}\rVert_{2}^{2}]$ after one step in the setting in which we are at the EoSS. Precisely we are computing here the value of $\mathbb{E}_{t}\mathbb{E}_{i}[\lVert Y_{i}^{t+1}\rVert_{2}^{2}]$ so the average over the iterations of the update to the quantity $\mathcal{C}$ above. Precisely we here prove the following Proposition.
Proposition 5.
In the setting and notations of Appendix B.1 and B.4. Assume $\eta\leq 2/\lambda_{\max}(\mathcal{H})$. Then there exists an absolute constant $c>0$ such that when Batch Sharpness   $>2/\eta+c\eta$, then $\mathcal{C}$ increases in size exponentially and the trajectory diverges (is quadratically unstable, see Definition 2). Note that assumptions on $GNI$ are not necessary.
Proof.
In the proof we use the notations of Appendix B.
Step 1: One step on the gradient’s second moment
. Remind that the SGD iterate satisfy
$\theta_{t+1}\;=\;\theta_{t}-\eta\,Y_{j_{t}}(\theta_{t}),\quad i_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{D},$
and define a fresh, independent index $j$ used only for the outer expectation in $\mathcal{C}^{t+1}$. Because $j\perp i_{t}$ we may write
$Y_{i}(\theta_{t+1})\;=\;H_{i}\!\bigl{(}\theta_{t+1}-x_{i}\bigr{)}=Y_{i}(\theta_{t})\;-\;\eta\,H_{i}Y_{j_{t}}(\theta_{t}).$
Squaring, expanding, and averaging over $j$ gives
$\begin{split}\mathcal{C}^{t+1}&=\mathbb{E}_{i}\bigl{\|}Y_{i}(\theta_{t})-\eta H_{i}Y_{j_{t}}(\theta_{t})\bigr{\|}^{2}\\ &=\mathcal{C}^{t}-2\eta\,\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{i}(\theta_{t})^{\!\top}H_{i}Y_{j_{t}}(\theta_{t})\bigr{]}}_{\text{cross term}}+\eta^{2}\,\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{j_{t}}(\theta_{t})^{\!\top}H_{i}^{2}Y_{j_{t}}(\theta_{t})\bigr{]}}_{\text{variance term}}.\end{split}$ (30)
Step 2: Decoupling the indices.
Note that (20) in the proof of Lemma 5 establishes that
$2\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{i}(\theta_{t})^{\!\top}H_{i}Y_{j_{t}}(\theta_{t})\bigr{]}\;=\;\mathcal{A}-\mathcal{B}-\widetilde{\Delta}.$ (31)
This implies that we can rewrite
$\begin{split}\mathcal{C}^{t+1}&=\mathcal{C}^{t}-\eta\,(\mathcal{A}-\mathcal{B}-\widetilde{\Delta})+\eta^{2}\,(\text{variance term}).\end{split}$ (32)
Next note that if we are at the EoSS, then $\mathcal{A}\approx\tfrac{2}{\eta}(1+\delta)\mathcal{C}$ for some $\delta\in\mathbb{R}$. This implies that we can rewrite the term above as
$\begin{split}\mathcal{C}^{t+1}\;&\approx\;-(1+2\delta)\,\mathcal{C}^{t}\;+\;\underbrace{\eta\mathcal{B}+\eta\widetilde{\Delta}+\eta^{2}\,(\text{variance term})}_{\text{rest}}.\end{split}$ (33)
Let us know understand the size of the rest, the trajectory diverges if and only if:
[…]
Next note that by applying Jensen inequality to the term multiplied by $\eta^{2}$ we obtain that
$\sqrt{\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{j_{t}}(\theta_{t})^{\!\top}H_{i}^{2}Y_{j_{t}}(\theta_{t})\bigr{]}}_{\text{variance term}}\cdot\underbrace{\mathbb{E}_{i}\!\bigl{[}Y_{i}(\theta_{t})^{\!\top}Y_{i}(\theta_{t})\bigr{]}}_{\mathcal{C}}}\;\geq\;\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{j_{t}}(\theta_{t})^{\!\top}H_{i}\cdot Y_{i}(\theta_{t})\bigr{]}}_{\mathcal{D}}.$ (35)
Step 3: Final algebra.
Plugging this above, we obtain that the trajectory diverges when
[…]
Again applying (31) we obtain that this is equivalent to
[…]
Since $\eta\mathcal{A}=2(1+\delta)\mathcal{C}$, then $\eta^{2}\mathcal{A}^{2}=4(1+\delta)^{2}\mathcal{C}^{2}$ to asking
[…]
Furthermore, equivalent to asking
$\eta\mathcal{B}+\eta\widetilde{\Delta}-\frac{2(1+\delta)}{2}\eta\widetilde{\Delta}-\frac{2(1+\delta)}{2}\eta\mathcal{B}+\eta^{2}\frac{\mathcal{B}^{2}+\widetilde{\Delta}^{2}+2\mathcal{B}\widetilde{\Delta}}{4\mathcal{C}}\;>\;(1-\delta+\delta^{2})\,\mathcal{C}$ (39)
or, even further simplified
$\eta\delta(\mathcal{B}+\widetilde{\Delta})+\eta^{2}\frac{(\mathcal{B}+\widetilde{\Delta})^{2}}{4\mathcal{C}}\;>\;(1-\delta+\delta^{2})\,\mathcal{C}.$ (40)
Here we plug in (20) again and we can rewrite this as
$\eta\delta(\mathcal{A}-2\mathcal{D})+\eta^{2}\frac{(\mathcal{A}-2\mathcal{D})^{2}}{4\mathcal{C}}\;>\;(1-\delta+\delta^{2})\,\mathcal{C}.$ (41)
By plugging, as before, $\eta\mathcal{A}=2(1+\delta)\mathcal{C}$ we obtain
$2\delta(1+\delta)\mathcal{C}-2\eta\delta 2\mathcal{D}-2\eta(1+\delta)\mathcal{D}+\eta^{2}\frac{\mathcal{D}^{2}}{\mathcal{C}}\;>\;\big{(}1-\delta+\delta^{2}-(1+\delta)^{2}\big{)}\,\mathcal{C}$ (42)
which simplifies as
$\underbrace{2\eta(1+2\delta)\mathcal{D}}_{\mathcal{O}(\eta^{2})}-\underbrace{\eta^{2}\frac{\mathcal{D}^{2}}{\mathcal{C}}}_{\mathcal{O}(\eta^{4})}\;<\;\delta(5+2\delta)\,\underbrace{\mathcal{C}}_{\mathcal{O}_{\eta}(1)}.$ (43)
Thus there exists a constant $c>0$, such that if $\delta>c\eta^{2}$ the trajectory diverges exponentially, if $\delta<c\eta^{2}$ the trajectory is stable.
Appendix D When $H_{i}\equiv H$: Pure Gradient–Noise Oscillations
Set–up.
Assume in the Setting of Appendix B every sample shares the same curvature: $H_{i}\equiv H\succeq 0$ for all $i$. SGD still sees noisy gradients $g_{t}=H\Delta_{t}+\xi_{t}$ with $\mathbb{E}_{i_{t}}[\xi_{t}]=0$, $\operatorname{Var}_{i_{t}}(\xi_{t})=\Sigma_{g}$, but now there is no curvature noise.
Batch Sharpness is capped by $\lambda_{\max}(H)$.
For any vector $v$, $v^{\top}Hv\leq\lambda_{\max}(H)\,\|v\|^{2}$. Averaging over the mini–batch therefore yields
$\textit{Batch Sharpness}(\theta)=\frac{\mathbb{E}_{i}[g^{\!\top}Hg]}{\mathbb{E}_{i}[\|g\|^{2}]}\;\leq\;\lambda_{\max}(H)\quad\forall\theta.$ (44)
Hence, as long as the classical stability condition $\eta<2/\lambda_{\max}(H)$ holds,
[…]
Gradient–Noise Interaction still plateaus at $\tfrac{2}{\eta}$.
The derivation in Appendix A did not use $H_{i}-H$ fluctuations, only the fixed stepsize. Consequently the stationary covariance satisfies $\Sigma_{x}=\eta\,\mathcal{K}^{\dagger}(\Sigma_{g})+\mathcal{O}(\eta^{2})$, and
$\mathrm{GNI}=\frac{\mathbb{E}_{i}[g^{\!\top}Hg]}{\|\nabla L\|^{2}}=\frac{2}{\eta}\bigl{[}1+\mathcal{O}(\eta)\bigr{]}.$ (45)
Thus SGD still wiggles with the universal Type–1 ratio $2/\eta$, even though the curvature never adapts.
Instability threshold reverts to $\lambda_{\max}(H)$.
Because (44) enforces Batch $\textit{Batch Sharpness}\leq\lambda_{\max}(H)$, the only way to reach the critical value $2/\eta$ is to push the stepsize past the quadratic limit:
[…]
Summary.
With $H_{i}\equiv H$ the dynamics exhibits only noise–driven (Type–1) oscillations:
  • $\mathrm{GNI}\!\to\!2/\eta$ at stationarity (same as the general case);
  • Batch Sharpness remains below $2/\eta$, bounded by $\lambda_{\max}(H)$;
  • the classical GD threshold $\eta=2/\lambda_{\max}(H)$ once again marks the onset of true instability.
In short, removing curvature variability collapses the Edge-of-Stochastic Stability back to the familiar quadratic picture.
Appendix E Mini-Batch Without Replacement
The main text and Appendix AC treated SGD with i.i.d. sampling. Here we show how the two key quantities
$\mathrm{GNI}\;=\;\frac{\mathbb{E}_{B}\!\bigl{[}g^{\!\top}\mathcal{H}g\bigr{]}}{\|\nabla L\|^{2}},\qquad\textit{Batch Sharpness}\;=\;\frac{\mathbb{E}_{B}[g^{\!\top}H_{B}g]}{\mathbb{E}_{B}[\|g\|^{2}]},$
behave when each epoch is a random permutation of the $n$ samples (batch size $b$, $k:=n/b$ steps per epoch).
We argue that without replacement sampling does not shift the EoSS in any practically relevant way. The only visible effect is a $1-\tfrac{b-1}{n-1}$ reduction in gradient-noise variance, which cancels in the GNI and Batch Sharpness and appears only as a second-order $\mathcal{O}(b/n)$ correction. The differences in effect between with and without replacements are thus about speed of convergence [49, 62] or in terms of traveling of the manifold of minima [36, 37], not in terms of EoSS.
Noise statistics under permutation sampling.
Fix $\theta$ and set $\bar{g}:=\nabla L(\theta)=\tfrac{1}{n}\sum_{i=1}^{n}\nabla\ell_{i}(\theta).$ For a single batch $B$ drawn without replacement
$g_{B}:=\frac{1}{b}\sum_{i\in B}\nabla\ell_{i}(\theta)=\bar{g}+\xi_{B},\qquad\mathbb{E}[\xi_{B}\mid\theta]=0,$
and the conditional covariance is [49, 25]
$\operatorname{Var}[\xi_{B}\mid\theta]=\frac{n-b}{b(n-1)}\;\Sigma_{g}(\theta),\quad\Sigma_{g}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\bigl{(}\nabla\ell_{i}-\bar{g}\bigr{)}\!\bigl{(}\nabla\ell_{i}-\bar{g}\bigr{)}^{\!\top}.$
Key point: compared to i.i.d. sampling the variance is simply scaled by $\alpha_{b,n}:=\tfrac{n-b}{n-1}\in(0,1].$
Updated Lyapunov solution.
Denote $\Delta_{t}:=\theta_{t}-\theta^{\star}$ and keep the stepsize fixed. The error recursion for a batch update is
$\Delta_{t+1}=(I-\eta\mathcal{H})\Delta_{t}-\eta\xi_{B_{t}},\qquad\xi_{B_{t}}\perp\!\!\!\perp\theta_{t}.$
Exactly as in Appendix A, the discrete Lyapunov equation now reads
$\Sigma_{x}=(I-\eta\mathcal{H})\Sigma_{x}(I-\eta\mathcal{H})+\eta^{2}\,\alpha_{b,n}\,\Sigma_{g}+\mathcal{O}(\eta^{3}),$
so that
$\boxed{\;\Sigma_{x}=\eta\,\alpha_{b,n}\,\mathcal{K}^{\dagger}(\Sigma_{g})+\mathcal{O}(\eta^{2}).\;}$
Correlation between $\xi_{B_{t}}$ and $\Delta_{t}$ enters only at $\mathcal{O}\!\bigl{(}\tfrac{b}{n}\eta\bigr{)}$ and is absorbed into the $\mathcal{O}(\eta^{2})$ remainder [see 37, App. E].
Gradient–Noise Interaction is mostly unchanged.
Plugging the equation above into the ratio $\mathbb{E}_{B}[g^{\!\top}\mathcal{H}g]/\|\nabla L\|^{2}$ shows that the common factor $\alpha_{b,n}$ cancels:
$\boxed{\;\mathrm{GNI}_{\text{w/o repl.}}=\frac{2}{\eta}\Bigl{[}1+\mathcal{O}(\eta)+\mathcal{O}\!\bigl{(}\tfrac{b}{n}\bigr{)}\Bigr{]}.\;}$
Hence the plateau $2/\eta$ persists up to tiny $\mathcal{O}(b/n)$ corrections.
Batch Sharpness and the instability edge.
  1. 1.
    Variable curvature ($H_{i}\!\not\equiv\!H$).  The numerator and the denominator of BatchSharpness contain the same prefactor $\alpha_{b,n}$, so their ratio—like $\mathrm{GNI}$—is unchanged at first order:
    $\textit{BatchSharpness}_{\text{w/o repl.}}=\frac{2}{\eta}\Bigl{[}1+\mathcal{O}(\eta)+\mathcal{O}\!\bigl{(}\tfrac{b}{n}\bigr{)}\Bigr{]}.$
    Consequently EoSS still appears once the directional curvature meets $2/\eta$; the $\mathcal{O}(b/n)$ shift is negligible whenever $n\gg b$.
  2. 2.
    Constant curvature ($H_{i}\equiv H$).  The bound $\textit{BatchSharpness}\leq\lambda_{\max}(H)$ of Appendix D is unaltered, so the classical quadratic threshold $\eta=2/\lambda_{\max}(H)$ remains the unique instability point.
Appendix F Hardware & Compute Requirements
All experiments were executed on a single NVIDIA A100 GPU (80 GB) with 256 GB of host RAM. The software stack comprises Python 3.12 and PyTorch 2.5.1 (built with the default CUDA tool-chain supplied by the wheel).
Baseline MLP (2M parameters, Section M)
Training for 200 epochs on the 8 k-image CIFAR-10 subset finishes in $\approx$ 5 min wall-clock while computing step sharpness every 8 updates and $\lambda_{\max}$ every 256 updates. Peak device memory is 14 GB during ordinary training and $\approx$ 70 GB while estimating $\lambda_{\max}$ on a 32k subset, comfortably fitting the 80 GB card.
Algorithmic caveats.
We rely on power iteration for $\lambda_{\max}$; while Lanczos would reduce the number of Hessian–vector products, the official PyTorch implementation remains CPU-only. To offset the extra memory incurred by double backward, we cache the first forward pass; batching $\lambda_{\max}$ is left to future work.
Appendix G Comparison with Previous Empirical Work
(a) Constant step size
(b) Step size increased early
(c) Step size increased late
Figure 11: We demonstrate that the saturation of GNI doesn’t govern a sharpness-related regime of instability typical of Type-2 oscillations - and in particular, highlighting the difference in the two types of oscillations. When we double the learning rate after batch sharpness is at least half of $2/\eta$ threshold (so that it is beyond the new $2/\eta$ level), training exhibits a catapult surge in the loss (c). But if we make the same change before batch sharpness crosses that level—despite GNI already saturating—no catapult occurs. (b)
(a) Constant batch size
(b) Batch size decreased early
(c) Batch size decreased late
Figure 12: Similarly, reducing the batch size only triggers catapults if batch sharpness, not GNI, exceeds the threshold.
Lee and Jang [17] introduce several quantities crucial for understanding neural network training dynamics. Below, we discuss the relationships among $\lambda_{\max}$, Batch Sharpness, and Interaction-Aware Sharpness (IAS, Lee and Jang [17]), emphasizing that a comprehensive theory of mini-batch dynamics should explain their distinct plateau timings and interconnected behaviors. We conjecture that a complete theory of stochastic gradient descent (SGD) dynamics would elucidate these metrics’ precise interrelations and their different plateau timings.
Interaction-Aware Sharpness
Lee and Jang [17] introduce Interaction-Aware Sharpness (IAS), denoted $\|\mathcal{H}\|_{S_{b}}$:
$\|\mathcal{H}\|_{S_{b}}\quad:=\quad\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\nabla L_{B}(\mathbf{x})^{\top}\mathcal{H}\,\nabla L_{B}(\mathbf{x})\big{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{B}\|^{2}\bigr{]}}.$
This quantity shares structural similarities with both Batch Sharpness (Definition 1) and the Gradient-Noise Interaction (Proposition 1), differing from the latter only in the denominator. The key distinction from Batch Sharpness lies in which Hessian is evaluated: IAS measures the directional curvature of the full-batch loss landscape $L$ along mini-batch gradient directions, while Batch Sharpness measures the directional curvature of mini-batch loss landscape $L_{B}$ along their corresponding gradients. This distinction is crucial, as mini-batch Hessians vary with batch selection while the full-batch Hessian remains fixed.
Notably, IAS serves as a directional alternative to the maximal Hessian eigenvalue, $\lambda_{\max}$, introduced by Cohen et al. [1]. IAS aligns closely with the $2/\eta$ threshold, unlike $\lambda_{\max}$, which often remains slightly above this threshold during EoS, especially at the beginning of it. Since IAS measures directional curvature, we have $\|\mathcal{H}\|_{S_{b}}\leq\lambda_{\max}$. Consequently, in the mini-batch setting, IAS stabilizes below $2/\eta$, consistent with empirical observations from Jastrzębski et al. [4, 5], Cohen et al. [1] and our Figure 2. Notably, when $B=n$, our Batch Sharpness coincides with IAS rather than $\lambda{\max}$, reinforcing the interpretation of Batch Sharpness as the relevant metric stabilizing at $2/\eta$ even under full-batch conditions.
Relation to Gradient-Noise Interaction
Another metric from Lee and Jang [17] is defined as:
$\frac{\operatorname{tr}(HS_{b})}{\operatorname{tr}(S_{n})}=\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\nabla L_{B}(\mathbf{x})^{\top}\mathcal{H}\,\nabla L_{B}(\mathbf{x})\big{]}}{\lVert\nabla L\rVert^{2}}$
which coincides exactly with our definition of GNI (Proposition 1). As detailed in Section 3 and Appendix A, the stabilization of GNI around $2/\eta$ signals the presence of oscillations, at least Type-1 oscillations. Lee and Jang [17] provide extensive empirical evidence demonstrating that neural networks spend much of their training within this oscillatory regime (see also Figures 11a and 12a). This contrasts traditional theoretical analyses (Bottou et al. [48], Mandt et al. [47]), which consider oscillations only near the manifold of minima.
Distinguishing oscillation types
It is crucial to note that GNI around $2/\eta$ does not inherently indicate instability. As clarified in Sections 3, 4 and Appendix A, not all oscillations are inherently unstable. Figures 4, 11b, 12b illustrate that altering hyperparameters when GNI is around $2/\eta$ typically does not trigger instability (catapult-like divergence), contrary to expectations if the system was in an EoS-like regime of instability. Instead, as shown in Figures 4, 11c, 12c, Batch Sharpness more reliably predicts a regime of instability. Additionally, Figure 13 highlights GNI’s independence from progressive sharpening, a necessary precursor to Type-2 (curvature-driven) oscillations and EoS-like instabilities, as detailed in Appendix A.5.
Missing Progressive Sharpening.
Extensively, both in our experiments and in the ones of Lee and Jang [17], GNI grows to $2/\eta$ in a few initial steps without ever being in subject to a phase of progressive sharpening unlike Batch Sharpness and $\lambda_{\max}$. The phase of growth of GNI is generally short and independent of the size, the behavior, and the phase in which Batch Sharpness and $\lambda_{\max}$ are.
[…]